# MonoPCC: Photometric-invariant Cycle Constraint for Monocular Depth Estimation of Endoscopic Images

### v1

@@ Abstract
Photometric constraint is indispensable for self-supervised monocular depth estimation. It involves warping a source image onto a target view using estimated depth&pose, and then minimizing the difference between the warped and target images. However, the endoscopic built-in light causes significant brightness fluctuations, and thus makes the photometric constraint unreliable. Previous efforts only mitigate this relying on extra models to calibrate image brightness. In this paper, we propose MonoPCC to address the brightness inconsistency radically by reshaping the photometric constraint into a cycle form. Instead of only warping the source image, MonoPCC constructs a closed loop consisting of two opposite forward-backward warping paths: from target to source and then back to target. Thus, the target image finally receives an image cycle-warped from itself, which naturally makes the constraint invariant to brightness changes. Moreover, MonoPCC transplants the source image’s phase-frequency into the intermediate warped image to avoid structure lost, and also stabilizes the training via an exponential moving average (EMA) strategy to avoid frequent changes in the forward warping. The comprehensive and extensive experimental results on three datasets demonstrate that our proposed MonoPCC shows a great robustness to the brightness inconsistency, and exceeds other state-of-the-arts by reducing the absolute relative error by at least 7.27%.
@@ 1 Introduction
Many efforts have been paid to enhance the reliability of photometric constraint under the fluctuated brightness. An intuitive solution is to calibrate the brightness in endoscopic video frames beforehand, using either a linear intensity transformation Ozyoruk et al. (2021) or a trained appearance flow model Shao et al. (2022). However, the former only addresses the global brightness inconsistency, and the latter increases the training difficulty due to introduced burdensome computation. Moreover, the reliability of appearance flow model is also not always guaranteed due to the weak self-supervision, which leads to a risk of wrongly modifying areas unrelated to brightness changes.
In this paper, we aim to address the bottleneck of brightness inconsistency without relying on any auxiliary model. Our motivation stems from a recent method named TC-Depth Ruhkamp et al. (2021), which introduced a cycle warping originally for solving the occlusion issue. TC-Depth warps a target image to source and then warps back to itself to identify every occluded pixel, since they assume the occluded pixel can not come back to the original position precisely. We find that such cycle warping can naturally overcome the brightness inconsistency, and yield a more reliable warped image compared to only warping from source to target, as shown in Fig. 1(b)-(d). However, directly applying the cycle warping often fails in photometric constraint because (1) the twice bilinear interpolation of cycle warping blurs the image too much, and (2) the networks of depth&pose estimation are learned actively, which makes the intermediate warping unstable and the convergence difficult.
@@
    Comprehensive and extensive experimental results on three public datasets demonstrate that MonoPCC outperforms other state-of-the-arts by decreasing the absolute relative error by at least 7.27%, and shows a strong ability to resist inconsistent brightness in the training.
@@ 2.1 Self-Supervised Depth Estimation
Self-supervised methods Zhou et al. (2017); Yin and Shi (2018); Bian et al. (2019); Godard et al. (2019); Watson et al. (2021) of monocular depth estimation mostly, if not all, rely on the photometric constraint to train two networks, named DepthNet and PoseNet. For example, as a pioneering work, SfMLearner Zhou et al. (2017) estimated depth maps and camera poses to synthesize a fake target image by warping another image from the source view, and calculated a reconstruction loss between the warped and target images as the photometric constraint.
For endoscopic cases, M3Depth Huang et al. (2022) enforced not only the photometric constraint but also a 3D geometric structural consistency by Mask-ICP. For improving pose estimation in laparoscopic procedures, Li et al. Li et al. (2022) designed a Siamese pose estimation scheme and also constructed dual-task consistency by additionally predicting scene coordinates. These methods share a common flaw: brightness fluctuations caused by the moving light source in the endoscope, significantly reduce the effectiveness of photometric constraint in self-supervised learning, degrading the performance consequently.
[…]
@@ 2.2 Brightness Inconsistency
Recently, several efforts have been made to mitigate brightness inconsistency in endoscopic cases, the common solution is to calibrate the brightness inconsistency using either preprocessing or extra learned models. For example, Endo-SfMLearner Ozyoruk et al. (2021) calculated the mean values and standard deviation of images, and then linearly aligned the brightness of the warped image with that of the target one. However, the real brightness fluctuations are actually local and non-linear, and are hardly calibrated by a simple linear transformation. Recently, AF-SfMLearner Shao et al. (2022) tried to relax the requirement of brightness consistency by learning an appearance flow model, which estimates pixel-wise brightness offset between adjacent frames. Nevertheless, effectiveness relies on the quality of the learned appearance flow model, and the risk of wrongly modifying the image content occurs accompanying the model error. Also, introducing an extra model dramatically increases the training time and difficulty in self-supervised learning.
@@ 3.1 Warping across Views for Self-Supervision
The DepthNet is typically an encoder-decoder, which inputs a single endoscopic image and outputs an aligned depth map. In this work, we use MonoViT Zhao et al. (2022) as our backbone DepthNet. The PoseNet contains a lightweight ResNet18-based He et al. (2016) encoder with four convolutional layers, which inputs a concatenated two images and predicts the pose change from the image in the tail channels to that in the front channels.
@@
where $p_{s}$ and $p_{s\rightarrow t}$ are the pixel’s homogeneous coordinates in $I_{s}$ and $I_{s\rightarrow t}$, respectively, $D(p)$ means the depth value of $D$ at the position $p$, and $K$ denotes the given camera intrinsic matrix. With the pixel matching relationship, $I_{s\rightarrow t}$ can be obtained by filling the color at each pixel position using differentiable bilinear sampling Jaderberg et al. (2015):
@@
where Structure Similarity Index Measure (SSIM) Wang et al. (2004) constrains the image quality and L1 loss constrains the content. $\alpha$ is set to 0.85 according to Godard et al. (2017).
One of the assumptions for the photometric constraint in Eq. (3) is that the scene contains no specular reflections, and is temporally consistent in terms of brightness Shao et al. (2022). However, such assumption hardly holds true in the endoscopic cases as visualized in Fig. 1.
@@ 3.3 Structure Transplant to Retain Details
The learning-free structure transplant module (STM) works based on two observations: (1) the warped frame $I_{t\rightarrow s}$ has homologous image style as $I_{t}$, and (2) the original frame $I_{s}$ has real structure details, and also is roughly aligned with $I_{t\rightarrow s}$. Inspired by the image-aligned style transformation proposed in Lv et al. (2023), we transplant fine structure of $I_{s}$ onto the appearance of $I_{t\rightarrow s}$ via a learning-free style transformation approach illustrated in Fig. 3.
Specifically, STM utilizes Fast Fourier Transform (FFT) Nussbaumer and Nussbaumer (1982) to decompose an RGB image $x$ into two components, i.e., amplitude $\mathcal{A}(x)$ and phase $\mathcal{P}(x)$. $\mathcal{A}(x)$ mainly controls the image style like brightness and $\mathcal{P}(x)$ contains the information of structure details Lv et al. (2023). Thus, after the forward warping, STM recombines $\mathcal{A}(I_{t\rightarrow s})$ and $\mathcal{P}(I_{s})$, and produces the structure-restored warped image $I_{t\rightarrow s}^{STM}$ via Inverse Fast Fourier Transform (IFFT):
@@ 3.5 Implementation Details
We train MonoPCC using a single NVIDIA RTX A6000 GPU. We utilize MonoViT and ResNet18 pre-trained on ImageNetDeng et al. (2009) for DepthNet and PoseNet initialization, respectively. We set the batch size to 12, and use AdamW Loshchilov and Hutter (2017) as the optimizer. The learning rate decay is exponential and set to 0.9. The number of epochs and learning rate differ in the warm-up and follow-up training phases. In warm-up training, the number of epochs is set to 20 epochs, and the initial learning rate is set to $1\times 10^{-4}$ for DepthNet’s encoder and $5\times 10^{-5}$ for the rest weights. In follow-up training, the number of epochs is set to 10 epochs, and the initial learning rate is set to $5\times 10^{-5}$ for all network weights.
@@ 4.1 Comparison Methods
We compare MonoPCC with 8 state-of-the-art (SOTA) methods of monocular depth estimation, i.e., Monodepth2 Godard et al. (2019), FeatDepth Shu et al. (2020), HR-Depth Lyu et al. (2021), DIFFNet Zhou et al. (2021), Endo-SfMLearner Ozyoruk et al. (2021), AF-SfMLearner Shao et al. (2022), MonoViT Zhao et al. (2022), and Lite-Mono Zhang et al. (2023). We evaluate their performance and make comparisons using their released codes.
4.2 Datasets
The experiment in this work involves 3 public endoscopic datasets, i.e., SCARED Allan et al. (2021), SimCol3D Rau et al. (2023), and SERV-CT Edwards et al. (2022).
Since the original data is binocular, we only use the left view to simulate our focusing monocular situation. We split the dataset into 21,066 frames for training and 551 for test, which are consistent with the previous SOTA Shao et al. (2022). No videos are cross-used in both training and test.
@@ 4.2 Datasets
Figure 5: The Abs Rel error maps of comparison methods on the three datasets. Opencv Jet Colormap is used for visualization.
@@ 4.3 Evaluation Metrics
To be consistent with previous works, we employ 5 metrics in the evaluation, which are listed in Table 1. Note that, as the common evaluation routine of monocular depth estimation, the predicted depth map should be scaled beforehand since it is scale unknown. The scaling factor is the GT median depth divided by the predicted median depth Zhou et al. (2017).
@@ 5.1.1 Evaluation on SCARED and SimCol3D
The top four rows of Fig. 5 present a qualitative comparison on the two datasets. Error maps are acquired by mapping pixel-wise Abs Rel value to different colors, and the red means large relative error and the blue indicates small error. On SCARED, the color-coded error map of MonoPCC is clearly bluer than those of other SOTAs, especially for the regions near the boundaries or specular reflections, as indicated by the red dotted boxes in Fig. 5.
On SimCol3D, MonoPCC provides lower relative depth errors at both the proximal and distal parts of the colon, which demonstrates the validity of MonoPCC for low-textured regions, e.g., colon in the digestive tract.
Methods Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ RMSE log $\downarrow$ $\delta$uparrow
Monodepth2 0.127 2.152 13.023 0.166 0.825
FeatDepth 0.117 1.862 12.040 0.154 0.841
HR-Depth 0.122 2.085 12.587 0.156 0.850
DIFFNet 0.116 1.858 12.177 0.146 0.864
Endo-SfMLearner 0.122 2.123 12.551 0.168 0.842
AF-SfMLearner 0.101 1.546 10.900 0.131 0.888
MonoViT 0.103 1.566 11.482 0.136 0.895
Lite-Mono 0.124 2.314 13.156 0.175 0.820
MonoPCC(Ours) 0.091 1.252 10.059 0.116 0.915
Table 3: Quantitative comparison results on SERV-CT. The best results are marked in bold and the second-best underlined. The paired p-values between MonoPCC and others are all less than $0.05$, except for $\delta$ compared to MonoViT.
5.1.2 Generalization on SERV-CT
We also evaluate the generalization of all methods, i.e., using the model trained on SCARED to directly test on SERV-CT. The comparison results are listed in Table 3. As can be seen, all methods experience a varying degree of degradation on SERV-CT. However, MonoPCC is still the best, and the only one maintaining the Abs Rel less than $0.1$ and $\delta$ greater than $0.9$. Also, MonoPCC significantly surpasses the second-best AF-SfMLearner by a relative decrease of 19.02% and 11.45% in the Sq Rel and RMSE log, respectively. The bottom two rows of Fig. 5 exhibit two visual results from SERV-CT. As can be seen, most methods yield higher errors compared to their visual results on SCARED. In comparison, MonoPCC still produces less red error maps as indicated by the dotted boxes in Fig. 5.
5.2 Ablation Study
@@
5.2.1 Effectiveness of Three Components
We develop five variants of MonoPCC by disabling the three components, EMA and/or $\mathcal{L}_{pcp}$, and STM. Note that, the warping should be cycle-form to utilize the components, and if we disable EMA, only $\theta$ in the backward warping path will be updated, and $\theta_{EMA}$ in the forward warping remains unchanged. Table 4 gives the comparison results between the variants and complete MonoPCC. We also include the backbone MonoViT in the first row of Table 4 as the baseline, which adopts the non-cycle warping.
From the comparison results in Table 4, three key observations can be made:
(1) By comparing the first two rows of Table 4, the inferior performance of the variant using none of the components indicates that the direct usage of cycle warping is not enough due to the image blurring and unstable gradient propagation mentioned in Sec. 3.2.
(2) As can be seen from the second to fourth rows, using EMA and $\mathcal{L}_{pcp}$ solely brings no improvements. However, adding STM can significantly reduce Sq Rel and RMSE by 16.11% and 7.50%, respectively ($p<0.003$). Thus, STM is the irreplaceable component for successfully training using the cycle-form warping.
(3) Comparison between the last four rows reveals the effectiveness of EMA and $\mathcal{L}_{pcp}$. Specifically, with STM enabled, both EMA and $\mathcal{L}_{pcp}$ bring remarkable reduction of Abs Rel by 8.08% ($p=0.012$, $p=0.022$). Meanwhile, EMA and $\mathcal{L}_{pcp}$ are not mutually excluded, and the combination reduces Abs Rel by 14.14% ($p=0.001$).

### v3

@@ Abstract
Photometric constraint is indispensable for self-supervised monocular depth estimation. It involves warping a source image onto a target view using estimated depth&pose, and then minimizing the difference between the warped and target images. However, the endoscopic built-in light causes significant brightness fluctuations, and thus makes the photometric constraint unreliable. Previous efforts only mitigate this relying on extra models to calibrate image brightness. In this paper, we propose MonoPCC to address the brightness inconsistency radically by reshaping the photometric constraint into a cycle form. Instead of only warping the source image, MonoPCC constructs a closed loop consisting of two opposite forward-backward warping paths: from target to source and then back to target. Thus, the target image finally receives an image cycle-warped from itself, which naturally makes the constraint invariant to brightness changes. Moreover, MonoPCC transplants the source image’s phase-frequency into the intermediate warped image to avoid structure lost, and also stabilizes the training via an exponential moving average (EMA) strategy to avoid frequent changes in the forward warping. The comprehensive and extensive experimental results on four endoscopic datasets demonstrate that our proposed MonoPCC shows a great robustness to the brightness inconsistency, and exceeds other state-of-the-arts by reducing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17%, respectively.
keywords:
Self-supervised learning , Monocular depth estimation , Photometric constraint , Brightness robustness
@@ 1 Introduction
Many efforts have been paid to enhance the reliability of photometric constraint under the fluctuated brightness. An intuitive solution is to calibrate the brightness in endoscopic video frames beforehand, using either a linear intensity transformation [19] or a trained appearance flow model [24]. However, the former only addresses the global brightness inconsistency, and the latter increases the training difficulty due to introduced burdensome computation. Moreover, the reliability of appearance flow model is also not always guaranteed due to the weak self-supervision, which leads to a risk of wrongly modifying areas unrelated to brightness changes.
In this paper, we aim to address the bottleneck of brightness inconsistency without relying on any auxiliary model. Our motivation stems from a recent method named TC-Depth [23], which introduced a cycle warping originally for solving the occlusion issue. TC-Depth warps a target image to source and then warps back to itself to identify every occluded pixel, since they assume the occluded pixel can not come back to the original position precisely. We find that such cycle warping can naturally overcome the brightness inconsistency, and yield a more reliable warped image compared to only warping from source to target, as shown in Fig. 1(b)-(d). However, directly applying the cycle warping often fails in photometric constraint because (1) the twice bilinear interpolation of cycle warping blurs the image too much, and (2) the networks of depth&pose estimation are learned actively, which makes the intermediate warping unstable and the convergence difficult.
@@
    We conduct comprehensive and extensive experiments on four public endoscopic datasets, i.e., SCARED [1], SimCol3D [21], SERV-CT [5], and Hamlyn [17, 26, 20] and a public natural dataset, i.e., KITTI [7]. The comparison results with eight state-of-the-art methods demonstrate the superiority of MonoPCC by decreasing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17% on the four endoscopic datasets, respectively, and its strong ability to resist inconsistent brightness in the training. Moreover, the comparison results on KITTI further verify the competitiveness of MonoPCC even for the natural scenario, where the brightness changes are often not that significant.
[…]
@@ 2.1 Self-Supervised Depth Estimation
Self-supervised methods [33, 29, 2, 9, 28] of monocular depth estimation mostly, if not all, rely on the photometric constraint to train two networks, named DepthNet and PoseNet. For example, as a pioneering work, SfMLearner [33] estimated depth maps and camera poses to synthesize a fake target image by warping another image from the source view, and calculated a reconstruction loss between the warped and target images as the photometric constraint.
For endoscopic cases, M3Depth [11] enforced not only the photometric constraint but also a 3D geometric structural consistency by Mask-ICP. For improving pose estimation in laparoscopic procedures, Li et al. [13] designed a Siamese pose estimation scheme and also constructed dual-task consistency by additionally predicting scene coordinates. These methods share a common flaw: brightness fluctuations caused by the moving light source in the endoscope, significantly reduce the effectiveness of photometric constraint in self-supervised learning, degrading the performance consequently.
@@ 2.2 Brightness Inconsistency
Recently, several efforts have been made to mitigate brightness inconsistency in endoscopic cases, the common solution is to calibrate the brightness inconsistency using either preprocessing or extra learned models. For example, Endo-SfMLearner [19] calculated the mean values and standard deviation of images, and then linearly aligned the brightness of the warped image with that of the target one. However, the real brightness fluctuations are actually local and non-linear, and are hardly calibrated by a simple linear transformation. Recently, AF-SfMLearner [24] tried to relax the requirement of brightness consistency by learning an appearance flow model, which estimates pixel-wise brightness offset between adjacent frames. Nevertheless, effectiveness relies on the quality of the learned appearance flow model, and the risk of wrongly modifying the image content occurs accompanying the model error. Also, introducing an extra model dramatically increases the training time and difficulty in self-supervised learning.
@@ 3.1 Warping across Views for Self-Supervision
The DepthNet is typically an encoder-decoder, which inputs a single endoscopic image and outputs an aligned depth map. In this work, we use MonoViT [31] as our backbone DepthNet. The PoseNet contains a lightweight ResNet18-based [10] encoder with four convolutional layers, which inputs a concatenated two images and predicts the pose change from the image in the tail channels to that in the front channels.
@@
where $p_{s}$ and $p_{s\rightarrow t}$ are the pixel’s homogeneous coordinates in $I_{s}$ and $I_{s\rightarrow t}$, respectively, $D(p)$ means the depth value of $D$ at the position $p$, and $K$ denotes the given camera intrinsic matrix. With the pixel matching relationship, $I_{s\rightarrow t}$ can be obtained by filling the color at each pixel position using differentiable bilinear sampling [12]:
@@
where Structure Similarity Index Measure (SSIM) [27] constrains the image quality and L1 loss constrains the content. $\alpha$ is set to 0.85 according to Godard et al. [8].
One of the assumptions for the photometric constraint in Eq. (3) is that the scene contains no specular reflections, and is temporally consistent in terms of brightness [24]. However, such assumption hardly holds true in the endoscopic cases as visualized in Fig. 1.
@@ 3.3 Structure Transplant to Retain Details
The learning-free structure transplant module (STM) works based on two observations: (1) the warped frame $I_{t\rightarrow s}$ has homologous image style as $I_{t}$, and (2) the original frame $I_{s}$ has real structure details, and also is roughly aligned with $I_{t\rightarrow s}$. Inspired by the image-aligned style transformation proposed in Lv et al. [15], we transplant fine structure of $I_{s}$ onto the appearance of $I_{t\rightarrow s}$ via a learning-free style transformation approach illustrated in Fig. 3.
Specifically, STM utilizes Fast Fourier Transform (FFT) [18] to decompose an RGB image $x$ into two components, i.e., amplitude $\mathcal{A}(x)$ and phase $\mathcal{P}(x)$. $\mathcal{A}(x)$ mainly controls the image style like brightness and $\mathcal{P}(x)$ contains the information of structure details [15]. Thus, after the forward warping, STM recombines $\mathcal{A}(I_{t\rightarrow s})$ and $\mathcal{P}(I_{s})$, and produces the structure-restored warped image $I_{t\rightarrow s}^{STM}$ via Inverse Fast Fourier Transform (IFFT):
@@ 3.5 Implementation Details
We train MonoPCC using a single NVIDIA RTX A6000 GPU. We utilize MonoViT and ResNet18 pre-trained on ImageNet[3] for DepthNet and PoseNet initialization, respectively. We set the batch size to 12, and use AdamW [14] as the optimizer. The learning rate decay is exponential and set to 0.9. The number of epochs and learning rate differ in the warm-up and follow-up training phases. In warm-up training, the number of epochs is set to 20 epochs, and the initial learning rate is set to $1\times 10^{-4}$ for DepthNet’s encoder and $5\times 10^{-5}$ for the rest weights. In follow-up training, the number of epochs is set to 10 epochs, and the initial learning rate is set to $5\times 10^{-5}$ for all network weights.
@@ 4.1 Comparison Methods
We compare MonoPCC with 8 state-of-the-art (SOTA) methods of monocular depth estimation, i.e., Monodepth2 [9], FeatDepth [25], HR-Depth [16], DIFFNet [32], Endo-SfMLearner [19], AF-SfMLearner [24], MonoViT [31], and Lite-Mono [30]. We evaluate their performance and make comparisons using their released codes.
@@
Figure 5: The Abs Rel error maps of comparison methods on SCARED and SimCol3D. Opencv Jet Colormap is used for visualization.
4.2 Datasets
The experiment in this work involves 4 public endoscopic datasets, i.e., SCARED [1], SimCol3D [21], SERV-CT [5], and Hamlyn [17, 26, 20], and a public outdoor-scene dataset, i.e., KITTI [7].
Since the original data is binocular, we only use the left view to simulate our focusing monocular situation. We split the dataset into 21,066 frames for training and 551 for test, which are consistent with the previous SOTA [24]. No videos are cross-used in both training and test.
Hamlyn is a large public endoscopic dataset with multiple stereo videos for clinical application, including roughly 8,000 frames with different resolutions. Recasens et al. [22] provided a rectified version and created GT depth maps for left-view frames using software Libelas.
KITTI is an outdoor dataset captured in autonomous driving scenarios, e.g., road and campus, while laser sensors are equipped to scan precise depth maps. We follow the split of Eigen et al. [6] and Zhou et al. [33], and use 39,810 frames for training, 4,424 for validation, and 697 for test.
@@ 4.3 Evaluation Metrics
To be consistent with previous works, we employ 5 metrics in the evaluation, which are listed in Table 1.
Note that, as the common evaluation routine of monocular depth estimation, the predicted depth map should be scaled beforehand since it is scale unknown. The scaling factor is the GT median depth divided by the predicted median depth [33].
Also, for a fair comparison, we follow the previous works [24, 21, 22, 9], and cap the depth maps within the maximum value for each dataset. Specifically, the maximum depth value is set to $150\ mm$, $200\ mm$, $180\ mm$, $300\ mm$, and $80\ m$ for SCARED, SimCol3D, SERV-CT, Hamlyn and KITTI, respectively.
Furthermore, for each comparison in the following experimental results, we perform paired samples T-test [4] and report the p-value to indicate whether the difference is significant.
@@ 5.1 Comparison with State-of-the-arts
Table 3: Quantitative comparison results on SERV-CT and Hamlyn. The best results are marked in bold and the second-best underlined. The paired p-values between MonoPCC and others are all less than $0.05$, except for $\delta$ on SERV-CT compared to MonoViT.
Methods SERV-CT Hamlyn
Monodepth2 0.127 2.152 13.023 0.166 0.825 0.087 1.254 9.555 0.115 0.939
FeatDepth 0.117 1.862 12.040 0.154 0.841 0.064 0.939 8.016 0.090 0.964
HR-Depth 0.122 2.085 12.587 0.156 0.850 0.076 1.139 8.883 0.102 0.957
DIFFNet 0.116 1.858 12.177 0.146 0.864 0.070 1.093 8.552 0.097 0.958
Endo-SfMLearner 0.122 2.123 12.551 0.168 0.842 0.063 0.886 7.901 0.091 0.968
AF-SfMLearner 0.101 1.546 10.900 0.131 0.888 0.078 1.018 8.712 0.104 0.968
MonoViT 0.103 1.566 11.482 0.136 0.895 0.073 0.945 8.470 0.099 0.975
Lite-Mono 0.124 2.314 13.156 0.175 0.820 0.074 1.106 8.743 0.104 0.950
MonoPCC(Ours) 0.091 1.252 10.059 0.116 0.915 0.061 0.819 7.604 0.085 0.977
Figure 6: The Abs Rel error maps of comparison methods on SERV-CT and Hamlyn. Opencv Jet Colormap is used for visualization.
@@ 5.1.1 Evaluation on SCARED and SimCol3D
Fig. 5 presents a qualitative comparison on the two datasets. Error maps are acquired by mapping pixel-wise Abs Rel value to different colors, and the red means large relative error and the blue indicates small error. On SCARED, the color-coded error map of MonoPCC is clearly bluer than those of other SOTAs, especially for the regions near the boundaries or specular reflections, as indicated by the red dotted boxes in Fig. 5. Specifically, the failure of other methods in these cases can be attributed to neglect or misguidance on brightness inconsistency, leading DepthNet to learn inappropriate depth prior knowledge. On SimCol3D, MonoPCC provides lower relative depth errors at both the proximal and distal parts of the colon, which demonstrates the validity of MonoPCC for low-textured regions, e.g., colon in the digestive tract.
@@
Figure 7: The Abs Rel error maps of seven ablation variants, including effectiveness of three components. (a)-(g) correspond to the $1^{st}$ to the $7^{th}$ rows in Table 4. Opencv Jet Colormap is used for visualization.
5.1.2 Generalization on SERV-CT and Hamlyn
We also evaluate the generalization of all methods, i.e., using the model trained on SCARED to directly test on SERV-CT and Hamlyn. The comparison results are listed in Table 3.
As can be seen, all methods experience a varying degree of degradation on both datasets. However, on SERV-CT (left part of Table 3), MonoPCC is still the best, and the only one maintaining the Abs Rel less than $0.1$ and $\delta$ greater than $0.9$. Also, MonoPCC significantly surpasses the second-best AF-SfMLearner by a relative decrease of 19.02% and 11.45% in terms of the Sq Rel and RMSE log ($p<0.005$), respectively.
Compared to the results on SERV-CT, the performance degradation on Hamlyn is somewhat not that severe. We believe that this is because the captured scenes of Hamlyn are relatively homogeneous, without complex and irregular surfaces as SERV-CT. Compared to the second-best Endo-SfMLearner, MonoPCC achieves the lowest error in terms of the first four metrics, and the highest percentage of the inliers in terms of $\delta$ ($p<0.001$ for all metrics).
Fig. 6 exhibits four visual results from SERV-CT and Hamlyn, respectively. As can be seen, compared to the other methods, MonoPCC produces less red error maps as indicated by the dotted boxes in Fig. 6. The two cases from Hamlyn contain motion blur during in vivo porcine procedure (the $3^{rd}$ row), and brightness fluctuations in the over-smoothed areas of silicon heart (the $4^{th}$ row). The color-coded error maps on the two challenging cases verify that MonoPCC displays almost deep blue error maps in the highlighted regions, and has better generalization on the non-rigid and reflective tissues.
5.2 Ablation Study